Circular No. 7876

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 C1 (IKEYA-ZHANG) H. E. Matthews, National Research Council, Victoria, British Colum-

bia, and Joint Astronomy Centre, Hawaii (JAC); and T. B. Lowe, JAC, report pointed spectral-line observations of comet C/2002 C1 made at the James Clerk Maxwell Telescope, Mauna Kea, in the 1.3-mm and 0.8-mm bands: “The beamwidths to halfpower are ∼ 21′′ and 13′′, respectively. We have detected the groundstate rotational transitions HCN 4–3 and 3–2, HNC 4–3 and 3–2 and CS 5–4. Observed integrated intensities (in mainbeam brightness K km/s) were: 3.65 and 0.71 (HCN 3–2, HNC 4–3, Mar. 23.95 UT), 4.10 (HCN 3–2, Mar. 28.82), 0.71 (HNC 3–2, Mar. 29.85), 18.4 (HCN 4–3, Mar. 30.98) and 4.47, 0.62 and 0.74 (HCN 3–2, HNC 3–2, CS 5–4, Apr. 7.93). CO has not been detected in two attempts (CO 3–2, Mar. 23.95 and CO 2–1, Apr. 7.93), the absolute upper limit to the peak signal being 0.1 K. The velocity width (km/s) to halfpower measured using the HCN lines was 2.6 (Mar. 23.93), 2.4 (Mar. 28.82), 2.2 (Mar. 30.98) and 2.2 (Apr. 7.87) and thus appears to show a decrease with increasing r.”

Ephemeris extension to IAUC 7858 (elements MPEC 2002-G38):

2002TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 16 23^{h}19.^{m}98 +55^{o}50.^{′}0 0.450 0.816 53.^{o}0 100.^{o}9 4.4
18 22 54.48 +57 48.3 0.438 0.849 57.0  97.4 4.5
20 22 25.33 +59 27.0 0.428 0.882 61.2  93.6 4.6
22 21 52.64 +60 40.1 0.419 0.916 65.7  89.7 4.7
24 21 17.05 +61 21.5 0.413 0.950 70.3  85.5 4.9
26 20 39.78 +61 26.5 0.408 0.984 75.1  81.3 5.0
28 20 02.50 +60 52.6 0.405 1.018 80.0  77.0 5.1
30 19 26.88 +59 41.1 0.404 1.051 84.9  72.6 5.2
May 2 18 54.19 +57 56.0 0.406 1.085 89.8  68.2 5.4
4 18 25.09 +55 43.4 0.410 1.119 94.7  63.9 5.6
6 17 59.70 +53 10.0 0.416 1.153 99.5  59.7 5.7

XTE J1751–305 Amplifying the private communication mentioned in the item by Wang

et al. on IAUC 7874, C. B. Markwardt, University of Maryland and Goddard Space Flight Center; and A. Dobrzycki, Harvard-Smithsonian Center for Astrophysics, report that on Apr. 10.67 UT the Chandra X-ray Observatory observed XTE J1751–305 with the HRC-S instrument. The source was detected at a level of 4.66 counts/s at the position α = 17^{h}51^{m}13^{s}.52, δ = −30^{o}37′22′′.9 (equinox 2000.0, 0′′.6 uncertainty).

2002 April 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden